OpenStudy (anonymous):

the change in x = vi x the change in t + 1/2a times the change in t squared. solve for a.

Parth (parthkohli):

\[\Delta x = vi \times \Delta t + {1 \over 2}a \times \Delta t^2\]The question doesn't make sense. Differentiation?

OpenStudy (anonymous):

modern phyics algebra review s - solving for a variable You have in typed correctly if you take out the multiplication signs.

OpenStudy (anonymous):

Can you tell me how you managed to type in the difference or delta symbol? That would help me type the problem in better. Thanks.

OpenStudy (anonymous):

I think he's simply saying: \[ \Delta x= v_i\Delta t+\frac{1}{2}a(\Delta t)^2 \]And you can use the equation editor or use brackets to express in \(\LaTeX\) whatever you wish.

OpenStudy (anonymous):

I don't know what the equation editor is.

OpenStudy (anonymous):

Solving for \(a\) is: \[\begin{align} \Delta x-v_i\Delta t&= \frac{1}{2}a(\Delta t)^2\implies\\ 2\Delta x-2v_i\Delta t&= a(\Delta t)^2\implies\\ \frac{2\Delta x-2v_i\Delta t}{(\Delta t)^2}&=a \end{align} \]

OpenStudy (anonymous):

The equation editor is at the bottom left-hand corner of the textbox for the answers.

OpenStudy (anonymous):

vf = vi + a times the change in t. solve for vi.